Ch 6.5 Dr. Mary Mandlebaum is a PAR provider in Medicare and in Mountville Health Plan, which has allowed charges for services and does not permit balance billing of plan members. She is not a PAR provider in the Ringdale Medical Plan. Based on the follow


A. Insurance Plan: Mountville Health Plan; patient has met annual deductible of $250; 80-20 coinsurance


Services: CPT 99203, 90700

Payer Reimbursement: ____________ Patient Charge: ____________




B. Insurance Plan: Mountville Health Plan; patient has paid $125 toward an annual deductible of $500; 80-20 coinsurance


Services: CPT 99215, 93040, 94010

Payer Reimbursement: ____________ Patient Charge: ____________




C. Insurance Plan: Ringdale Medical Plan A; no deductible or coinsurance; copayment of $5/PAR; $25/NonPAR


Services: CPT 99212

Payer Reimbursement: ____________ Patient Charge: ____________




D. Insurance Plan: Ringdale Medical Plan B; patient has met annual deductible of $300; 80-20 coinsurance


Services: CPT 99215

Payer Reimbursement: ____________ Patient Charge: ____________




E. Insurance Plan: Medicare; annual deductible has been met by patient


Services: 99213, 93040

Payer Reimbursement: ____________ Patient Charge: ____________
Answer #1

A. Mountville Health Plan’s allowed charges are the basis for billing.

Payer: Allowed Charges: $89 + $87 = $176 × 80% = $140.80;

Patient: $176 × 20% = $35.20

B. Mountville Health Plan’s allowed charges are the basis for billing.

Payer: Allowed Charges: $119 + $36 + $69 = $224.

Patient has not met $375 of the deductible, so patient pays entire charge of $224.

C. The provider’s usual charges are the basis for billing. The patient pays a $25 copayment. Payer pays $44 – $25 = $19.

D. The provider’s usual charges are the basis for billing. The payer pays $134 x 80% = $107.20; patient $134 × 20% = $26.80.

E. The Medicare Allowed Charge is the basis for billing.

Payer: Allowed charges: $48 + $30 = $78 × 80% = $62.40;

Patient: $78 × 20% = $15.60